OpenStudy (anonymous):

Can anyone please help? A chessboard has an area of 324 sq. inches. There is a 1-inch border around the 64 squares on the board. What is the length of one side of the region containing the small squares?

OpenStudy (mathmate):

 Since x^2=324, we get x=sqrt(324)=18 => x-2=18-2=16 A chess board has 8x8=64 squares. Can you work out the rest?